MIPO_15_10_13

MIPO_15_10_13
CW, AG, PG, MP, AN, PMH, AB, PS, -­‐ KL Actions SG action on magnet crash buttons-­‐ complete. FNAL have document, still under review. Some issues have been identified. MP assembling detailed reply, expect end of week. FC status • FC is cold DL currently checking power supply. • Running -­‐if there are ramps to quench, should monitor quench on SS. • Should run into short first to fully characterise power supplies. • Agree to review results of those tests into short before proceeding. • circulate test results to MIPO. SS status • AB, technical update, ran full current last wednesday • magnet was very good. diode trays have water leaks -­‐ now fixed • SSU looks good 1e-­‐8 vac 3-­‐4 watts extra cooling capacity. • SSD M1 coil not operable • have good model, no short term fix, looking into work past step IV. • M2 test also completed at 5A nothing anomalous. Considered operational. • Pick-­‐up on M1 read through voltage taps becomes a qd for m2. • Finished design of DCCT system including buffers to do digital calculation and protect m2. • Plan to run SSD if we have step IV running. • Evaluate power supply system, after magnet review. • Hope to have DCCT completed the week of magnet review. • Ground fault detection will monitor short between M1 and M2. AB sequence of events. • internal qp blew a lead on M1, not clear why, due to delay in quench -­‐ longer current drive into QP? • When centre coil quenches, centre coil qp does absorb current for 25 seconds-­‐ longer than time to damage in fault event. Normally M1 would only see current for 10s. • could be intrinsic design problem or single point failure -­‐ have checked photographs taken during build • qp opens, voltage rises, circuit is open, results in arc, guess at feedthroughs from vac to he space • That burnt out lead in M1 and caused short between M1 and M2, break is in LTS lead. • centre coil has 2 diodes vs 1 in match solenoids. SSU&D are currently not being monitored on advice of AB. • MP we left with statement magnets should be monitored, were hoping isis shifters to check periodically and call expert. • AB changed this advice as vacuum and He were very stable. PMH no change in past few days. • JB plan would be to monitor vac and heater power and he level. PS should we monitor T? AB probably not necessary. Magnet power supplies. • Should SSD dump trays be refitted? Yes refit. Allows running pre-­‐final step IV config. • AB Centre coil on tray was 125C during run, maybe add fans? has been noted by SG. • FC ground fault and quench protection installed, test today at low current TH. • MP full configurational safety review. Reviewing suggestions from SG and analysis, expect some changes. • Need to validate multi-­‐magnet operation. How power supply is configured at system level. QP integration is needed. • MP notes components are in place and need to be validated. Need a meeting of 1 day from experts. MP need a lead for detailed systems integration review. after MPB/RLSR. • Magnet Review Oct 26th. • Booked CR13 R68. 1PM, Jim Kerby at DoE request, initial assessment. • More detailed review planned for nov23-­‐24th at FNAL. • DoE recommends to MPB that after the technical review is complete and we have a recovery plan we should have a sub-­‐committee of MPB and RLSR review the plans. Which they may choose to modify. • MP notes members have been unable to contact Charlotte. • MP RAL to confirm room for magnet review. Might be better to ask for CR12. • Action on CW to change to CR12, provide coffee etc. RLSR/MPB -­‐ CW/KL Uk finance meeting today. in good shape US to meet today to finalise risk table and updated spend profile, expect to get send to Uk in next 24 hours. Safety • Incident has precipitated a protest. • DS power supply out of hours intervention, nobody at RAL was advised. • AN recommends we need full time run co-­‐ordinator. CW hope PH, V.Blackmore, V.Bayliss, C.McWaters can combine to cover this. • AN meeting with Neil Geddes later this week. • AN meeting with divison head, AO, expect critcism. • AN lab and experiment operate under an HSE framework with safety codes -­‐ there is electrical safety code which details power supply interventions -­‐ need electric permit by authorised person, which did not happen last wed when ds power supply tripped due to over temp. •
•
•
•
This was a reportable occurrence, Report of incident is accessible -­‐ SHE enterprise online PRY drawings -­‐ CW AB last week verified that JT and SP are working from same cooling optics document. • CW why do we have a requirement for the RF cavity to be able to be turned through 180, AB -­‐ in case we have to put in a spare. Related to services. • Steve Plate to communicate with AB by end of week and AB to circulate. Need to resolve before end next week. • If we decide to use MTA magnet we would require a new pry. RF Plan -­‐ AG Slides posted Applied recommendations from RF review to increase time contingency for risks showing additional effort as well as ASTEC effort and Bob Anderson -­‐ ISIS. Result is • pushed system 1 delivery back. • system 2 is also pushed back. • moving start date forward will not help as rf is limiting factor. moving back will not immediately impact critical path. • high power testing finishes beginning of July, then PRY can be installed. • AG picked out critical path and risk added critical path. • Risk might change. Schedule reflects best understanding of RF only. KL has asked for a meeting with CJ, BS, IR, etc before review to advise of current status Hydrogen system update. • hydrogen system hardware is complete • He cooldown is in preparation -­‐ 1 flexible pipe to be installed and safety paperwork to be done. • Next action is preliminary tour once feedback has been incorporated, should be ready to run. AOB. 
Was this manual useful for you? yes no
Thank you for your participation!

* Your assessment is very important for improving the work of artificial intelligence, which forms the content of this project

Download PDF

advertising